OpenStudy (anonymous):

Solve. 2x^2 + x = 15.

OpenStudy (anonymous):

use -b formula, where a=2 b=1 and c=-15,

OpenStudy (radar):

Factor (x+3)(2x-5)=0

OpenStudy (anonymous):

2x^2 + x - 15 = 0 2x^2 +3x - 5x - 15 = 0 2x(x +3 ) - 5(x + 3) = 0 (2x - 5 )(x + 3) = 0

OpenStudy (anonymous):

x = 5/2 . -3

OpenStudy (anonymous):

x = 5/2 and x = -3

OpenStudy (anonymous):